# Test-time Adaptive Hierarchical Co-enhanced Denoising Network for Reliable Multimodal Classification

### v1

@@ IV Experiments
TABLE I: Comparison of classification performance between the proposed TAHCD and other reliable multimodal classification methods on the BRCA, ROSMAP, CUB, and UPMC FOOD101 datasets, where both training and testing are conducted under the same noise settings. The best results under various noise settings are indicated in bold, while the second-best results are underlined.
@@
TAHCD 84.4 84.0 80.0 81.1 81.1 85.6 92.8 92.8 92.6 93.0 93.1 93.1
$\eta=10\%$ $\epsilon=0$ MD [12] 70.8 69.7 54.5 72.6 73.9 78.1 69.4 69.3 68.3 73.1 73.2 73.2
MLCLNet [49] 70.4 68.2 50.5 70.4 70.6 76.2 69.2 68.8 68.1 73.2 73.9 74.0
QMF [47] 68.8 64.9 50.3 70.8 69.3 73.1 70.3 70.2 69.5 73.4 74.3 74.3
PDF [3] 65.0 56.7 41.7 67.9 72.8 77.3 69.4 69.3 69.0 74.1 74.2 74.2
NCR [18] 69.6 65.4 51.1 69.8 70.9 76.3 65.1 65.3 65.3 72.1 72.3 72.2
ALBEF [22] 69.1 65.2 50.8 70.0 73.1 77.6 65.8 65.7 65.7 72.4 72.3 72.3
@@
TAHCD 73.4 73.0 69.4 79.2 79.6 83.5 73.0 73.3 72.6 75.5 75.6 75.6
$\eta=10\%$ $\epsilon=5$ MD [12] 65.8 58.5 42.4 68.9 72.7 72.3 68.5 67.9 67.2 73.3 73.4 73.4
MLCLNet [49] 65.1 58.2 42.1 64.9 70.1 71.5 67.6 67.7 68.3 72.9 72.8 73.1
QMF [47] 65.0 58.9 44.4 62.2 69.2 70.1 64.8 65.1 64.8 73.2 73.7 73.7
PDF [3] 60.5 48.1 31.0 66.0 72.7 72.1 68.5 68.0 67.8 72.2 72.9 72.9
NCR [18] 64.2 65.0 45.2 64.7 70.2 68.0 67.1 67.1 67.9 72.0 72.8 72.8
ALBEF [22] 67.9 66.0 48.3 68.3 71.1 70.8 69.4 69.6 69.4 72.3 72.4 72.2
@@
SPS [43] 68.1 66.2 48.3 69.2 72.4 71.3 70.2 70.1 70.9 72.8 72.8 72.5
TAHCD 69.2 67.9 60.1 76.4 78.6 77.2 73.0 73.0 72.4 75.3 75.1 75.2
TABLE II: Comparison of classification performance between the proposed TAHCD and other reliable multimodal classification methods on the BRCA, ROSMAP, CUB, and UPMC FOOD101 datasets. All methods are trained on the original datasets and test on data corrupted with multimodal noise $(\eta=10\%,\epsilon=5)$.
@@
IV-A Experimental Settings
IV-A1 Datasets
Experiments are conducted on four commonly used multimodal datasets in previous works [13, 11, 12, 47, 49, 50, 4]. (1) BRCA: BRCA [28] is a dataset for breast invasive carcinoma PAM50 subtype classification. The dataset comprises 875 samples, with each sample containing features from three modalities: mRNA expression data (mRNA), DNA methylation data (meth), and miRNA expression data (miRNA). These samples are categorized into five subtypes: Normal-like, Basal-like, HER2-enriched, Luminal A, and Luminal B, with 115, 131, 46, 435, and 147 samples, respectively. BCRA can be obtained from The Cancer Genome Atlas program (TCGA)  [^1] https://www.cancer.gov/aboutnci/organization/ccg/research/structuralgenomics/tcga. (2) ROSMAP: ROSMAP [30, 1, 6] is a dataset containing samples from Alzheimer’s patients and normal control subjects. The dataset consists of 351 samples, including 182 Alzheimer’s disease patients and 169 normal control samples. Each sample includes data from three modalities: mRNA expression data (mRNA), DNA methylation data (meth), and miRNA expression data (miRNA). (3) CUB: Caltech-UCSD Birds dataset [39] comprises 200 categories of birds. It contains a total of 11,788 samples, with each sample including data from two modalities: images of birds and their corresponding textual descriptions. (4) UPMC FOOD101: The UPMC FOOD101 dataset [41] comprises food images from 101 categories obtained through Google Image search and corresponding textual descriptions. This dataset contains 90,704 samples, where each sample’s image and text are collected from uncontrolled environments, thus inherently containing noise.
IV-A2 Compared Methods
To validate the reliability of TAHCD under multimodal noise, we compare it with several representative reliable multimodal learning methods in our experiments, including multimodal dynamics (MD) [12], multi-level confidence learning (MLCLNet) [49], quality-aware multimodal fusion (QMF) [47], predictive dynamic fusion (PDF) [3], Noisy Correspondence Rectifier (NCR) [18], Align before Fuse (ALBEF) [22], Semantic Invariance Learning (SMILE) [44], and Seeking Proxy Point via Stable Feature Space (SPS) [43]. The first four methods focus on learning high-confidence modality representations or fusing different modalities based on their qualities, and are therefore regarded as representative approaches for modality-specific noise removal. The remaining four methods emphasize learning cross-modality consistency and correspondence, and are thus treated as representative approaches for cross-modality noise removal.
TABLE III: Ablation studies on four datasets with noise ($\eta=10\%,\epsilon=5$) added on both training and test data
@@ IV-A2 Compared Methods
72.7 73.8 74.0
53.1 52.7 53.8
IV-A3 Evaluation Metrics
BRCA & CUB & UMPC FOOD101
ROSMAP
IV-A4 Implementation Details
Training Details
Model Implementation Details
To evaluate the classification performance of NCR [18] and ALBEF [22], the sample features after being filtered or rectified are utilized to train the classifier. For SMILE [44] and SPS [43], the sample features output by its well-trained encoders are used for classification.
Experimental Details
Two types of noise are involved in the experiments: modality-specific and cross-modality noise. The modality-specific noise is implemented following previous works [11, 50, 3], and $\epsilon$ is its severity level. The implementation of cross-modality noise follows SMILE [44], where unaligned samples are introduced by shuffling the features of each modality for a random subset of samples with a proportion of $\eta$.
TABLE IV: Ablation studies on four datasets with training on original data and testing under noise added data ($\eta=10\%,\epsilon=5$).
@@ IV-A4 Implementation Details
54.8 55.9 54.8
52.2 52.3 50.8
IV-B Comparison with State-of-the-Art Methods
To demonstrate the improved robustness and generalization ability of the proposed TAHCD under multimodal noise, we conduct comparative experiments on datasets corrupted by modality-specific and cross-modality noise, following previous works [12, 3, 44]. $\epsilon$ and $\eta$ represent the severity levels of modality-specific and cross-modality noise, respectively.
IV-B1 Robustness
To evaluate the improved robustness of TAHCD against multimodal noise, we compare the classification performance of different methods on training and test sets that are both corrupted by the same noise, as reported in Table I. Methods such as MD, MLCLNet, QMF, and PDF learn reliable modality representations by estimating modality confidence but lack awareness of cross-modality alignment. Consequently, their performance is relatively stable under settings with only modality-specific noise ($\eta=0,\epsilon=5$), while their performance degrades significantly once cross-modality noise is introduced (all settings with $\eta=10\%$). For NCR, ALBEF, SMILE, and SPS, these methods primarily focus on cross-modality alignment while paying limited attention to confidence of each modality. As a result, they exhibit relatively stable performance when only cross-modality noise is present ($\eta=10\%,\epsilon=0$), but their performance degrades noticeably once modality-specific noise is introduced (all settings with $\epsilon=5$). In contrast, TAHCD consistently achieves more stable performance than the compared methods under various noise conditions on all datasets, demonstrating superior robustness. This advantage primarily stems from its explicit modeling of both modality-specific and cross-modality noise and their joint removal across multiple levels, enabling more thorough noise suppression.
IV-B2 Generalization ability
To evaluate the improved generalization ability of TAHCD under previously unseen noise, we train all methods on the original datasets and test them on noise-corrupted data. Table II reports the test results on data corrupted by a mixture of modality-specific noise with $\epsilon=5$ and cross-modality noise with $\eta=10\%$. When confronted with noise unseen during training, all baseline methods suffer substantial performance degradation. In particular, compared with the results under the noise setting of $\eta=10\%$ and $\epsilon=5$ in Table I, their performance degrades substantially, indicating limited adaptability to previously unseen noise during training. In contrast, our TAHCD outperforms other baselines when exposed to previously unseen noise, while exhibiting a noticeably smaller performance degradation compared to the setting where such noise is observed during training. The possible reason is that TAHCD can adapt in response to input noise at test time, thereby exhibiting stronger generalization ability.
IV-C Abliation Study
@@
To evaluate the overall effectiveness of the three key components of TAHCD, namely Adaptive Stable Subspace Alignment (ASSA), Sample-Adaptive Confidence Alignment (SACA), and Test-Time Cooperative Enhancement (TTCE), we conduct ablation studies. As shown in Table III, following the same setting as Table I, identical noise is added to both the training and test data to assess the contribution of each component to model robustness. The results indicate that all three components contribute to robustness improvement, with ASSA providing the largest gain, followed by SACA and TTCE. A possible reason is that the noise is sampled from a shared distribution and added to all samples, exhibiting strong global characteristics, which allows ASSA to remove most of the noise at the global level. SACA also contributes to performance improvements, indicating that even noise sampled from the same distribution contains outlier components, thereby demonstrating the necessity of instance-level denoising on each sample. The performance gains brought by TTCE indicate that the cooperative enhancement between global-level and instance-level denoising is beneficial for improving denoising effectiveness.
@@
Additionally, to evaluate the contribution of each component to generalization ability, we follow the same setting as in Table II by training the model on the original dataset and testing it on noise-corrupted data. The results are shown in Table IV. Removing TTCE leads to a substantial performance drop, indicating its critical role in adapting to previously unseen noise and improving generalization. When TTCE is removed, further removing ASSA or SACA results in additional performance degradation, indicating that these two components also contribute to learning more generalizable multimodal representations. A possible reason is that ASSA and SACA learn informative modality distribution priors, which help guide the model to better remove unseen noise.
IV-D Discussion On ASSA
As detailed in Sec. III-B, ASSA learns a mask $w_{\lambda}^{m}$ based on the singular values obtained from the decomposition of modality feature covariance matrices to construct stable feature subspace, guided by two constraints $\mathcal{L}_{\mathrm{o}}$ and $\mathcal{L}_{\mathrm{a}}$.
IV-D1 Discussion On Adaptive Stable Subspace Construction
We evaluate the effectiveness of stable subspace construction by examining whether ASSA learns noise-related masks $w^{m}_{\lambda}$. Specifically, we generate Gaussian noise $n^{m}\in\mathbb{R}^{N\times d^{m}}$ with severity $\epsilon=5$, set the values of half of the feature dimensions in the noise matrix to zero, and add the resulting noise to one modality $x^{m}$. We then visualize the learned mask $w^{m}_{\lambda}$ before and after noise corruption in Fig. 3, with results shown for the corrupted image modality on the CUB dataset. Before adding noise (left panel of Fig. 3), the learned mask show small values in few feature dimensions, suggesting that most features of the original high-quality data are informative and largely retained. After the noise is added (right panel of Fig. 3), approximately half of the learned mask values corresponding to corrupted feature dimensions decrease, appearing lighter in the figure. This indicates that ASSA can learn masks highly correlated with noise, thereby preserving informative principal axes and ensuring the stability of the constructed feature subspace.
@@ IV-D1 Discussion On Adaptive Stable Subspace Construction
TABLE V: Evaluation of the effects of $\mathcal{L}_{\mathrm{o}}$ and $\mathcal{L}_{\mathrm{a}}$ on datasets with modality-specific noise ($\epsilon=5$).
@@
TABLE VI: Evaluation of the effects of $\mathcal{L}_{\mathrm{o}}$ and $\mathcal{L}_{\mathrm{a}}$ on datasets with cross-modality noise ($\eta=10\%$).
@@
Proposed 79.2 79.6 83.5
TABLE VII: Effectiveness of Alignment on Projections onto Stable Feature Subspaces under Cross-Modality Noise with Varying $\eta$ (with Modality-Specific Noise $\epsilon=5$).
@@
BRCA w/o projection 84.1 65.2 54.8
with projection 85.6 73.4 65.8
ROSMAP w/o projection 85.6 74.1 60.3
with projection 87.7 79.2 71.8
CUB w/o projection 91.5 68.8 57.9
with projection 92.8 73.0 66.2
FOOD101 w/o projection 92.0 70.2 59.6
with projection 93.2 75.5 68.4
IV-D2 Effectiveness of $\mathcal{L}_{\mathrm{o}}$ and $\mathcal{L}_{\mathrm{a}}$
Further experiments are conducted to evaluate the roles of the two constraints $\mathcal{L}_{\mathrm{o}}$ and $\mathcal{L}_{\mathrm{a}}$ in noise removal. We perform ablation studies on datasets corrupted only by modality-specific noise with $\epsilon=5$ and only by cross-modality noise with $\eta=10\%$, respectively. As shown in Table V, on datasets with only modality-specific noise $(\epsilon=5)$, removing $\mathcal{L}_{\mathrm{o}}$ causes a much more significant performance drop than removing $\mathcal{L}_{\mathrm{a}}$. This indicates that $\mathcal{L}_{\mathrm{o}}$ plays a dominant role in guiding modality-specific noise removal. In contrast, as shown in Table VI, on datasets with only cross-modality noise $(\eta=10\%)$, removing $\mathcal{L}_{\mathrm{a}}$ results in a substantially larger performance drop compared to removing $\mathcal{L}_{\mathrm{o}}$, demonstrating that $\mathcal{L}_{\mathrm{a}}$ primarily guides cross-modality noise removal.
IV-D3 Further Discussion On the Subspace Projection Alignment
@@
To further verify the effectiveness of Subspace Projection Alignment in aligning projections onto stable feature subspaces, we add cross-modality noise with varying intensity $\eta$ on data with modality-specific noise $\epsilon=5$. We compare two strategies: aligning directly on modality features (w/o projection) and aligning projections onto stable subspaces via Subspace Projection Alignment (with projection). As shown in Table VII, under modality-specific noise, the “with projection” strategy exhibits stronger robustness than “w/o projection” as cross-modality noise increases. This indicates that aligning modality projections within their respective stable subspaces enables more effective suppression of cross-modality noise in the presence of modality-specific corruption.

### v2

@@ 4.2.1 Robustness
Table 1: Comparison of classification performance between the proposed TAHCD and other reliable multimodal classification methods on the BRCA, ROSMAP, CUB, and UPMC FOOD101 datasets, where both training and testing are conducted under the same noise settings. The best results under various noise settings are indicated in bold, while the second-best results are underlined.
@@
ASL [32] 75.5 73.2 69.1 73.4 74.1 81.5 85.1 85.0 85.2 91.1 91.5 91.5
DisNCL [9] 76.3 73.4 69.8 74.2 74.6 81.9 85.2 85.3 85.3 91.2 91.5 91.4
TAHCD 84.4 84.0 80.0 81.1 81.1 85.6 93.6 93.5 93.5 94.0 94.1 94.1
$\eta=10\%$ $\epsilon=0$ MD [13] 65.0 56.7 41.7 67.9 72.8 77.3 69.4 69.3 69.0 73.1 73.2 73.2
MLCLNet [39] 70.4 68.2 50.5 70.4 70.6 76.2 69.2 68.8 68.1 73.2 73.0 73.1
QMF [37] 68.8 64.9 50.3 70.8 69.3 73.1 70.3 70.2 69.5 74.4 74.3 74.3
PDF [4] 70.8 69.7 54.5 72.6 73.9 78.1 69.4 69.3 68.3 73.1 73.2 73.2
@@
ASL [32] 71.9 71.4 62.8 74.7 76.2 82.3 71.7 71.6 71.6 74.2 74.3 74.3
DisNCL [9] 72.0 72.3 62.6 75.1 76.3 82.4 71.9 71.8 72.0 74.8 74.6 74.5
TAHCD 77.4 77.0 69.4 79.2 79.6 84.5 76.4 76.3 76.6 77.5 77.6 77.6
$\eta=10\%$ $\epsilon=5$ MD [13] 60.5 48.1 31.0 66.0 72.7 72.1 68.5 68.0 67.8 69.2 69.3 69.3
MLCLNet [39] 65.1 58.2 42.1 64.9 70.1 71.5 67.6 67.7 68.3 69.9 69.8 70.1
QMF [37] 65.0 58.9 44.4 62.2 69.2 70.1 64.8 65.1 64.8 71.2 71.7 71.7
PDF [4] 65.8 58.5 42.4 68.9 72.7 72.3 68.5 67.9 67.2 72.3 72.4 72.4
@@
SPS [33] 68.1 66.2 48.3 69.2 72.4 71.3 70.2 70.1 70.9 71.8 71.8 71.5
ASL [32] 68.2 66.5 50.2 70.4 73.1 71.8 70.4 70.5 70.5 72.0 71.8 71.7
DisNCL [9] 68.5 67.1 50.9 70.9 73.3 72.5 70.8 70.9 70.9 72.2 72.3 72.3
TAHCD 73.2 71.9 60.1 76.4 78.6 77.2 75.3 75.2 75.4 76.3 76.1 76.2
4.2.2 Generalization ability
To evaluate the improved generalization ability of TAHCD under previously unseen noise, we train all methods on the original datasets and test them on noise-corrupted data. Table 2 reports the results on test data corrupted by a mixture of modality-specific noise with $\epsilon=5$ and cross-modality noise with $\eta=10\%$. When confronted with noise unseen during training, all baseline methods suffer substantial performance degradation. Specifically, compared with the results under the noise setting of $\eta=10\%$ and $\epsilon=5$ in Table 1 (where noise is added on both train and test sets), their performance in Table 2 degrades substantially, indicating limited adaptability to previously unseen noise during training. In contrast, our TAHCD outperforms other baselines when exposed to previously unseen noise, while exhibiting a noticeably smaller performance degradation compared to the setting where such noise is observed during training. The possible reason is that TAHCD can better adapt in response to input noise at test time, thereby exhibiting stronger generalization ability.
Table 2: Comparison of classification performance between the proposed TAHCD and other reliable multimodal classification methods on the BRCA, ROSMAP, CUB, and UPMC FOOD101 datasets. All methods are trained on the original datasets and test on data corrupted with multimodal noise $(\eta=10\%,\epsilon=5)$.
@@ 4.2.2 Generalization ability
Table 3: Ablation studies on four datasets under noise ($\eta=10\%,\epsilon=5$).
(a) Noise added on both training and testing data.
@@
76.4 78.6 77.2
(b) Noise only added on testing data.
@@
75.1 77.3 76.4
4.3 Abliation Study
@@
To evaluate the overall effectiveness of the three key components of TAHCD, namely Adaptive Stable Subspace Alignment (ASSA), Sample-Adaptive Confidence Alignment (SACA), and Test-Time Cooperative Enhancement (TTCE), we conduct ablation studies. As shown in Table 3(a), following the same setting as Table 1, identical noise is added to both the training and test data to assess the contribution of each component to model robustness. The results indicate that all three components contribute to robustness improvement, with ASSA providing the largest gain, followed by SACA and TTCE. A possible reason is that the noise is sampled from a certain distribution and added to all samples, exhibiting strong global characteristics, which allows ASSA to remove most of the noise at the global level. SACA also contributes to performance improvements, indicating that even noise sampled from the same distribution contains outlier components, thereby demonstrating the necessity of instance-level denoising on each sample. The performance gains brought by TTCE indicate that the cooperative enhancement between global-level and instance-level denoising is beneficial for improving robustness.
@@
Additionally, to evaluate the contribution of each component to generalization ability, we follow the same setting as in Table 2 by training the model on the original dataset and testing it on noise-corrupted data. The results are shown in Table 3(b). Removing TTCE leads to a substantial performance drop, indicating its critical role in adapting to previously unseen noise and improving generalization. When TTCE is removed, further removing ASSA or SACA results in additional performance degradation, indicating that these two components also contribute to learning more generalizable multimodal representations. A possible reason is that ASSA and SACA learn informative modality distribution priors, which help guide the model to better remove unseen noise.
@@
4.4 Discussion On ASSA
As detailed in Sec. 3.2, ASSA learns a mask $w_{\lambda}^{m}$ based on the singular values obtained from the decomposition of modality feature covariance matrices to construct stable feature subspace, guided by two constraints $\mathcal{L}_{\mathrm{o}}$ and $\mathcal{L}_{\mathrm{a}}$.
4.4.1 Discussion On Adaptive Stable Subspace Construction
We evaluate the effectiveness of stable subspace construction by examining whether ASSA learns noise-related masks $w^{m}_{\lambda}$. Specifically, we generate Gaussian noise $n^{m}\in\mathbb{R}^{N\times d^{m}}$ with severity $\epsilon=5$, set the values of half of the feature dimensions in the noise matrix to zero, and add the resulting noise to one modality $x^{m}$. We then visualize the learned mask $w^{m}_{\lambda}$ before and after noise corruption in Figure 3, with results shown for the corrupted image modality on the CUB dataset. Before adding noise (left panel of Figure 3), the learned mask show small values in few feature dimensions, suggesting that most features of the original high-quality data are informative and largely retained. After the noise is added (right panel of Figure 3), approximately half of the learned mask values corresponding to corrupted feature dimensions decrease, appearing lighter in the figure. This indicates that ASSA can learn masks highly correlated with noise, thereby preserving informative principal axes and ensuring the stability of the constructed feature subspace.
Table 4: Evaluation of the effects of $\mathcal{L}_{\mathrm{o}}$ and $\mathcal{L}_{\mathrm{a}}$ on datasets under noise.
(a) Under modality-specific noise ($\epsilon=5$).
@@ 4.4.1 Discussion On Adaptive Stable Subspace Construction
(b) Under cross-modality noise ($\eta=10\%$).
@@
Proposed 79.2 79.6 84.5
4.4.2 Effectiveness of $\mathcal{L}_{\mathrm{o}}$ and $\mathcal{L}_{\mathrm{a}}$
Further experiments are conducted to evaluate the roles of the two constraints $\mathcal{L}_{\mathrm{o}}$ and $\mathcal{L}_{\mathrm{a}}$ in noise removal. We perform ablation studies on datasets corrupted only by modality-specific noise with $\epsilon=5$ and only by cross-modality noise with $\eta=10\%$, respectively. As shown in Table 4(a), on datasets with only modality-specific noise $(\epsilon=5)$, removing $\mathcal{L}_{\mathrm{o}}$ causes a much more significant performance drop than removing $\mathcal{L}_{\mathrm{a}}$. This indicates that $\mathcal{L}_{\mathrm{o}}$ plays a dominant role in guiding modality-specific noise removal. In contrast, as shown in Table 4(b), on datasets with only cross-modality noise $(\eta=10\%)$, removing $\mathcal{L}_{\mathrm{a}}$ results in a substantially larger performance drop compared to removing $\mathcal{L}_{\mathrm{o}}$, demonstrating that $\mathcal{L}_{\mathrm{a}}$ primarily guides cross-modality noise removal.
Table 5: Accuracy of alignment on projection and without projection under cross-modality noise with varying $\eta$ (with modality-specific noise $\epsilon=5$).
@@ 4.4.2 Effectiveness of $\mathcal{L}_{\mathrm{o}}$ and $\mathcal{L}_{\mathrm{a}}$
BRCA w/o projection 84.1 60.2 51.8
with projection 84.4 73.2 72.8
ROSMAP w/o projection 80.7 71.2 58.3
with projection 81.1 76.4 71.8
CUB w/o projection 93.1 67.8 57.9
with projection 93.6 75.3 70.2
FOOD101 w/o projection 93.8 69.2 58.6
with projection 94.0 76.3 70.4
4.4.3 Further Discussion On the Subspace Projection Alignment
@@
To further verify the effectiveness of Subspace Projection Alignment in aligning projections onto stable feature subspaces, we add cross-modality noise with varying intensity $\eta$ on data with modality-specific noise $\epsilon=5$. We compare two strategies: aligning directly on modality features (w/o projection) and aligning projections onto stable subspaces via Subspace Projection Alignment (with projection). “w/o projection” is achieved by directly minimizing the Frobenius norm of differences between modality features. As shown in Table 5, under modality-specific noise, the “with projection” strategy exhibits stronger robustness than “w/o projection” as cross-modality noise increases. This indicates that aligning modality projections within their respective stable subspaces enables more effective suppression of cross-modality noise in the presence of modality-specific noise.
4.5 Discussion On SACA
4.5.1 Effectiveness of Noise Experts
More experiments are conducted to verify that the modality specific noise experts and cross modality noise experts in SACA respectively facilitate the removal of instance level modality-specific and cross-modality noise. Specifically, we first introduce only modality-specific noise $\epsilon=5$ into one modality of the dataset and visualize the sample masks $\{w_{s,i}^{m}\}_{i=1}^{N}$ learned by the corresponding modality-specific noise expert. Figure 4 illustrates the masks learned by the modality-specific noise expert on the image modality affected by noise in the CUB dataset. In the noise-free setting (left), the mask values exhibit relatively strong consistency across samples. After modality-specific noise is added (right), the learned mask shows increased variability between samples, showing that the modality-specific noise expert effectively captures the instance-level chaos introduced by the modality noise. Also, we introduce cross-modality noise with $\eta=20\%$ to one modality of the dataset by randomly shuffling the features of 20% of its samples. Figure 5 shows the masks learned by the cross-modality noise expert on the image modality affected by noise in the CUB dataset. Before noise is added (left), the masks of different samples exhibit a relatively consistent pattern, whereas after noise introduction (right), the masks show substantial variability across samples. This demonstrates that the cross-modality noise expert can also effectively capture sample-specific corrupted information introduced by noise.
Table 6: Comparison of model accuracy under noise using different cross-modality alignment methods.
Dataset Method $\eta=0$ $\eta=10\%$ $\eta=20\%$
BRCA sim 84.6 66.8 57.3
MI 85.0 70.8 61.3
slack 85.6 77.4 72.8
ROSMAP sim 86.2 75.3 62.8
MI 86.9 76.4 65.1
slack 87.7 79.2 71.8
CUB sim 92.0 69.8 60.5
MI 92.3 70.6 62.7
slack 93.8 76.4 70.2
FOOD101 sim 92.5 71.3 61.8
MI 93.0 72.6 64.1
slack 94.2 77.5 70.4
4.5.2 Discussion on Confidence-Aware Asymmetric Slack Alignment
We conduct experiments to validate the effectiveness of the slack alignment proposed in Confidence-Aware Asymmetric Slack Alignment. Specifically, we compare with existing methods that enforce cross-modality consistency, including similarity-based approaches (sim), as used in [33, 9], and information-theoretic approaches (MI), as used in [34]. “slack” represents applying our $\mathcal{L}_{\mathrm{nll}}^{c}$ that retains the slack-alignment mechanism while removing modality-confidence guidance. As shown in Table 6, under cross-modality noise with varying intensity $\eta$, our slack alignment exhibits stronger stability than other methods, indicating superior robustness. “MI” achieves better performance than “sim”, likely due to its less restrictive constraint on cross-modality consistency, preserving more complementary information across modalities.
We also conduct experiment to validate the effectiveness of incorporating modality confidence to guide slack alignment. Specifically, the model is trained on data with modality-specific and cross-modality noise ($\epsilon=1,\eta=10\%$), and the changes in modality-wise confidence and the loss $\mathcal{L}_{\mathrm{nll}}^{c}$ are reported for both with and without confidence guidance. The variant without confidence guidance is implemented by removing the gradient modulation in Eq. (13). Modality confidence reported in Figure 6 is computed as the average of the likelihoods $p^{m}(\hat{h}_{s,i}^{m})$ of all sample representations $\hat{h}_{s,i}^{m}$ under the corresponding modality-specific feature prior $p^{m}$. As shown in the left panel of Figure 6, as training progresses, the confidence of each modality gradually increases, indicating progressive noise removal. Moreover, with modality-confidence guidance, modality confidence rises to higher levels than without such guidance, suggesting that confidence guidance promotes alignment toward high-confidence features across modalities, thereby yielding more robust representations. Moreover, as shown in the right panel of Figure 6, the loss $\mathcal{L}_{\mathrm{nll}}^{c}$ decreases faster with confidence guidance than without it, indicating that confidence guidance also accelerates the alignment process.
Figure 6: Changes in confidence of each modality and the decline of $\mathcal{L}_{\mathrm{nll}}^{c}$ during training on the FOOD101 dataset with cross-modality noise $(\eta=10\%)$ and modality-specific noise $(\epsilon=1)$, under two conditions: with and without the guidance of modality confidence on slack alignment.